OpenStudy (anonymous):

Rewrite the expression 10cos^4(x) as an equivalent expression that does not contain the power of trigonometric functions greater than 1.

OpenStudy (luigi0210):

Use the idenitity \(\Large cos^2x=\frac{1-cos(2x)}{2}\)

OpenStudy (luigi0210):

*+\(cos2x\) not -

OpenStudy (thomas5267):

Use \cos for cosine. It is nicer.

OpenStudy (anonymous):

This is the answer I got:\[(5/2)-5\cos2x+(5/2)\cos^22x\] I know it's wrong. Where did I go wrong?

OpenStudy (anonymous):

oooh I see now

OpenStudy (anonymous):

nope. I'm still getting a wrong answer

OpenStudy (anonymous):

@Luigi0210 @thomas5267

OpenStudy (thomas5267):

Use \frac{a}{b} for \(\frac{a}{b}\). Simplify the \(\dfrac{5}{2}\cos^2(2x)\) again using \(\cos^2(\theta)=\dfrac{\cos(2\theta)+1}{2}\)

OpenStudy (anonymous):

I got it! Thank you @thomas5267 and @luigi0210